OpenStudy (anonymous):

Find the simple interest. principal = $750; rate = 14%; time = 4.5 years A. $105 B. $420 C. $472.50 D. $1222.50

pooja195 (pooja195):



pooja195 (pooja195):

Start by turning the percentage into a decimal

OpenStudy (anonymous):

0.01

pooja195 (pooja195):

To conver 14% into a decimal you divide it by 100 14/100=?

OpenStudy (anonymous):

0.14

pooja195 (pooja195):

Ok good now we plug it into the formula \[\huge~\rm~P=750 \times 0.14 \times 4.5\] calculate that :)

pooja195 (pooja195):

Let me know when you have the answer

OpenStudy (anonymous):

P = 472.5

pooja195 (pooja195):

Correct which option does that match ?

pooja195 (pooja195):

P=472.50 dont forget the 0 :P

OpenStudy (anonymous):

C!

pooja195 (pooja195):

Yes! ^_^ Well done! :-)